OpenStudy (anonymous):

Write an equation of a parabola with the given focus and directrix focus: (0,3) directrix: y=-1

OpenStudy (ikram002p):



OpenStudy (ikram002p):

ok now could you tell me how to write parabola equation with focus and directrix ?

OpenStudy (anonymous):

\[y=a(x-h)^{2}+k\]

OpenStudy (anonymous):

that's the equation for a parabola, right? @ikram002p

OpenStudy (ikram002p):

and what is k what is h ?

OpenStudy (anonymous):

h=0 k=3 ?

OpenStudy (ikram002p):

a parabola with the a focus (a,b ) and directrix y=c is givven by (x-a)^2+b^2-c^2=2(b-c)y

OpenStudy (ikram002p):

so u have a=0 b=3 c=-1 write the equation :)

OpenStudy (ikram002p):

it will be (x-0)^2 + 3^2- (-1)^2 =2(3+1) y